MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2007/02/16
OBSERVER: Kamiya Paul
CONDITION: Clear
COMMENT: No good darks. Because chip 2, 3 & 4 are noisy. Problem is fixed, but sky is very bright.
Compressor gas pressure: 2.2 MPa
Expected Observation Time: START 22:05 END 06:00
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 4857-4876 | 20 | 87 | 5220 |
LMC | 1225-1233 | 9 | 50 | 15000 |
SMC | 330-330 | 1 | 2 | 600 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 30 | 139 | 20820 (05h47m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
9 | 7 | 5 | 4 | 15 | 3 | 4 | 3 | 11 | 3 | 3 | 2 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
2 | 2 | 2 | 2 | 2 | 2 | 2 | 2 | 1 | 1 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
2 | 2 | 2 | 2 | 2 | 2 | 9 | 9 | 9 | 3 | 2 | 2 |
Field13 | Field14 | Field15 | Field16 |
2 | 2 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CHECK SHEET
Time | Status | Weather |
22:00 | observing | clear |
23:00 | observing | clear |
0:00 | observing | clear |
1:00 | observing | clear |
2:00 | observing | clear |
3:00 | observing | clear |
4:00 | observing | clear |
5:00 | observing | clear |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day